# Exploring the Boundaries of On-Device Inference: When Tiny Falls Short, Go Hierarchical

### v1

@@ 4.2.1. Image transmission: WiFi vs BLE
As expected, BLE’s power consumption is smaller, $2.23$ W, compared to WiFi, which consumes $2.71$ W. However, the transmission time of an image under BLE is $177.13$ ms, two orders of magnitude higher than the $2.65$ ms time it takes for WiFi. Thus, the average energy consumption for transmitting an image to the server and receiving two bytes back consumes 7.39 mJ in the case of Wi-fi and 395 mJ in the case of BLE. Given significantly high latency and energy consumption under BLE, we opt for Wi-Fi communication for offloading the images.
4.2.2. Measuring offloading time over WiFi
@@
We measure the total time and on-device energy consumption for transmitting an image over Wi-Fi and getting the inference back at the device. The device, operating as a client, transmits a request to the server that is actively monitoring communications. To establish Wi-Fi connectivity, the client connects to a hotspot that access the local network where the server is operating. The experiment was designed to transmit an image as a raw byte array via TCP and receive a number representing the anticipated class by performing inference using the remote ML on the server using the incoming array. The duration between the initiation of the offload process and the reception of the response is repeated and averaged over $10000$ images. Measurements for this process with both datasets are presented in Table 3.
4.3. Device Results
@@
The results for base models were acquired on 3 devices (Coral Micro, Raspberry Pi, Jetson Orin) are shown in Table 4. The results for Coral Micro illustrate that for CIFAR-10, AlexNet is inferior to ResNet models in all performance metrics. For ImageNet-1K, the AlexNet model we trained, cannot be loaded on the device due to its size, even when modifying the board’s memory regions via the linker script, in order to allocate more space for it. For the Raspberry Pi, all the models from Subsection 3.2 were successfully run on the device. For Jetson Orin, an important observation is that operating at $15$W results in lower energy consumption than $7$W power. This is because, at high power, the latency is lower. This observation is akin to BLE vs WiFi offloading time.
@@
Regarding CIFAR-10, Jetson Orin is the most efficient in terms of both energy and latency when compared to the other two devices. However, for Raspberry Pi, the tradeoff is the power efficiency, which is not as efficient as Coral Micro for ResNet-8 ($\times$2.9) and ResNet-56 ($\times$5.6). However, one anomaly is Alexnet, which works significantly more efficient on Raspberry Pi ($\times$5.8) than on Coral Micro.
@@
For ImageNet-1K, a similar pattern is observed. Yet, when it comes to latency, Raspberry Pi surpasses Coral Micro for ResNet-50 by a factor of 1.5. This indicates that running more complex models on Coral Micro results in additional processing time, even with the assistance of a TPU accelerator.
4.4. On-Device vs. HI vs. Remote Inference
@@
In Fig. 3, we present the accuracy achieved by different strategies for CIFAR-10 and ImageNet-1K datasets, respectively. We present the latency and energy consumption of these strategies on Coral Micro, Raspberry Pi, and Jetson Orin in Figs. 4 and Fig. 5, respectively. ResNet-8 and ResNet-18 are the SOTA tinyML models for CIFAR-10 and ImageNet-1K, respectively, and are used as the local DL models for the HI system. From Figs. 4, and 5, we observe similar trends across the CIFAR-10 and ImageNet-1K datasets for both Coral Micro and Raspberry Pi. In all the subfigures, the smallest local ML models, ResNet-8 for CIFAR-10 and ResNet-18 for ImageNet-1K, consume the least energy and have the lowest latency among the models. However, these models also have the lowest accuracies (cf. Fig. 3) and do not satisfy the accuracy QoS requirements.
a Accuracy on CIFAR-10
b Accuracy on ImageNet-1K
Figure 3. Accuracy of different strategies on CIFAR-10 and ImageNet-1K datasets
a Latency for CIFAR-10
b Energy consumption for CIFAR-10
@@
Figure 4. Latency and energy comparison of on-device and HI systems for CIFAR-10 on different devices.
a Latency for ImageNet-1K
b Energy consumption for ImageNet-1K
@@
Figure 5. Latency and energy comparison of on-device and HI systems for ImageNet-1K on different devices.
@@
The HI system provides better accuracy—up to $2.2$% higher for CIFAR-10 (cf. Fig. 3a) and $11.5$% higher for ImageNet-1K (cf. Fig. 3b) than both on-device models. This comes with only a slight increase in latency and energy compared to the smallest model (cf. Figs. 4 and 5). Remote ML inference, while yielding the highest accuracy, for CIFAR-10 results in significantly higher latency and energy consumption on both Coral Micro and Raspberry Pi, as shown in Figs. 4a and 4b. However, for ImageNet-1K, offloading the data for remote ML inference results in lower energy consumption and latency due to the increase in the local ML model size and dataset complexity (cf. Figs. 5a, and 5b). In contrast, Jetson Orin, with its GPU acceleration, is highly suitable for on-device inference, as larger models like ResNet-56 for CIFAR-10 and ResNet-50 for ImageNet consume significantly less latency and energy compared to remote ML inference (cf. Figs. 4, and 5).
@@
Next, we compare the performances of three different systems: on-device inference, HI system, and remote inference, based on the QoS requirements defined in Section 3.3. For Arduino Nano and ESP32, in Fig. 1 we already showed that doing on-device inference using LR meets the accuracy QoS for the MNIST dataset, with much lower latencies than offloading. However, for CIFAR-10, the SOTA tinyML ResNet-8 fails to meet the $90$% QoS requirement, necessitating a larger model. However, high latencies for ResNet-8 on these MCUs suggest even higher latency and energy for larger models, making remote inference the preferred strategy to meet QoS requirements.
@@
For Coral Micro and Raspberry Pi, we observe from Figs. 4a and 4b that, for CIFAR-10 dataset, the on-device ResNet-56 model and the HI system with ResNet-8 can achieve the required accuracy QoS of $90$%. Although HI systems add some latency and energy overheads to the on-device model ResNet-8, these are insignificant compared to the much higher latency and energy required for ResNet-56 on-device inference. Consequently, the HI system meets the accuracy QoS with up to $73$% lower latency and $77$% lesser energy than the on-device model. Given the latency and energy QoS requirements (cf. Table 3) on these two devices, the HI system achieved more than a $2$% absolute accuracy gain. For Jetson Orin, on-device inference using ResNet-56 is much faster and more energy-efficient than the HI system. Thus, for the $90$% accuracy requirement, on-device inference suffices. Since the HI system achieves a higher accuracy ($1.64$% more) while still satisfying the energy and time QoS requirements, HI is the best strategy in those scenarios.
We observe from Table 4 and Fig. 3b that for ImageNet-1K dataset, both the on-device DL models, ResNet-18, and ResNet-50, with accuracies $69.7$% and $76$%, fall far behind from the required accuracy QoS of $80$%. Moreover, both these models demand significantly more energy and latency on Coral Micro and Raspberry Pi compared to remote inference (cf. Figs. 5a and 5b). This makes using even a larger local model in those two devices infeasible. On the other hand, Offloading data samples for remote inference achieves higher accuracy with much lower latency and energy. Hence, remote inference is the best strategy for the $80$% accuracy QoS requirement. Further, remote inference is also the fastest and most energy-efficient solution in this scenario.
On Jetson Orin, however, the on-device inference is faster and more energy-efficient. Therefore, we choose the RegNetY32GF model for on-device inference, as both ResNet-18 and ResNet-50 fail to meet the $80$% accuracy requirement. Using RegNetY32GF, we achieved $82.09$% accuracy with an inference time of $14.88$ms and energy consumption of $213.67$mJ, which is higher than that of remote inference. In contrast, the HI system with ResNet-18 meets the accuracy QoS with $50$% lower latency and $72$% lesser energy. Additionally, HI improves accuracy by $2.14$% over RegNetY32GF while satisfying both latency and energy QoS. The best strategies for each QoS requirement are summarized in Table 5.
Table 5. Preferred strategy for QoS requirements.
@@
It is noteworthy that the decision module is crucial to the success of HI systems as demonstrated in Eq. 2. In our experiments, binary LR performed adequately, achieving F1 scores of 0.86 on ResNet-8 for CIFAR-10 and 0.83 on ResNet-18 for the ImageNet-1K dataset. A key insight from the above performance comparison is that:
@@
  • In general, for resource-constrained devices, relatively complex tasks (CIFAR-10 for Arduino Nano and ESP32, ImageNet-1K for Coral Micro and Raspberry Pi) require much higher energy and latency to achieve QoS requirements with on-device inference. Consequently, offloading these tasks for remote inference is a better choice.
@@
  • If the on-device inference is feasible, it is more beneficial to opt for an HI system with a smaller model that doesn’t satisfy the QoS requirement than a larger on-device model that does. HI systems often achieve QoS requirements while optimizing latency and energy consumption (for accuracy QoS and vice versa).
5. Early Exit with HI
Figure 6. EE-HI for a Convolutional Neural Network (CNN) with a single EE branch.
Figure 7. EE, HI, EE-HI comparison for different models for CIFAR-10 on Raspberry Pi.
5.1. Algorithm Description
Equations 3 and 4 show that the latency and energy consumption of a given HI system is a function of four variables: (i) on-device performance, (ii) decision module, (iii) offloaded samples ratio, and (iv) offload impact, which includes both the communication cost and the computation cost at the server. Therefore, optimizing any of these variables can enhance the performance of an HI system.
According to inference strategies discussed so far, each sample must pass through the entire local ML model before being either accepted or offloaded. This implies that a fixed latency and energy overhead is added to the system. As existing techniques such as early exit (Teerapittayanon et al., 2016) have been shown to reduce on-device overhead while maintaining similar accuracy, we evaluate how introducing this strategy can enhance an HI system. As mentioned earlier, we refer to the case of early exit with HI as EE-HI.
The intuition behind early exit is to create a model capable of assessing how much computation is needed for a given sample. To achieve this, additional branches are attached to the model at different positions, each with its own exit point. During inference, a sample is first evaluated at the initial branch. If the confidence level is sufficient, the system accepts the output and proceeds with the next sample. Otherwise, inference continues to the subsequent branch. Thus, no extra overhead is added unless necessary.
An overview of what an EE-HI system looks like with CNNs is shown in Fig. 6. Several variables need to be considered when designing an early exit architecture, as they significantly impact the performance of the model. Some of the most important variables are the number, placement, and composition of the branches. Nonetheless, selecting these variables is mostly an empirical process, as stated in (Laskaridis et al., 2021).
@@ 5.1. Algorithm Description
The design and training of the models follows the pioneering BranchyNet approach (Teerapittayanon et al., 2016) and the training details can be found in Subsection 3.2. We place the branches using a trial-and-error method. For ResNet-8, we use one EE branch after the second residual block or fifth convolutional layer, and for AlexNet, we use one EE branch after the second convolutional layer. For ResNet-56, we use two EE branches, noting that larger models allow for more branches without adding excessive overhead relative to the original latency. The first branch is placed after layer $2$, and the second one after layer $19$.
@@
Table 6. Results for the EE-HI system with optimal threshold subject to the QoS requirement.
@@
5.2. Results
In order to compare the performance of this novel strategy, a constraint needs to be set on either accuracy, latency, or energy consumption. This allows us to choose the optimal combination of early exit thresholds that satisfy the given conditions. We adhere to the previously defined minimum accuracy QoS requirement for CIFAR-10, i.e., 90% . We find the optimal threshold combination for each model using a brute-force approach. These results are presented in Table 6, along with those obtained for HI. These experiments are conducted on Raspberry Pi and Jetson Orin using the CIFAR-10 dataset.

### v2

@@ V-B1 Image transmission – WiFi vs BLE
Given the availability of both BLE and Wi-Fi on all devices (except Arduino Nano), a key question arises: which one should be used? To explore this, we conducted an experiment using a Raspberry Pi, measuring both latency and energy consumption while transmitting a CIFAR-10 image to the server and waiting for a response. It is worth mentioning that the energy measurements account for the entire working device, not just the communication module. As expected, BLE’s power consumption is smaller, $2.23$ W, compared to WiFi, which consumes $2.71$ W. However, the transmission time of an image under BLE is $177.13$ ms, two orders of magnitude higher than the $2.65$ ms time it takes for WiFi. Thus, the average energy consumption for transmitting an image to the server and receiving two bytes back consumes 7.39 mJ in the case of Wi-Fi and 395 mJ in the case of BLE. Given the significantly high latency and energy consumption under BLE, we opt for Wi-Fi communication for offloading the images.
V-B2 Measuring offloading time over WiFi
@@
The total time and device energy consumption of offloading include the transmission of the image over Wi-Fi and receiving the inference back at the device. The device, operating as a client, sends a request to the server, which actively monitors communications. To establish Wi-Fi connectivity, the client connects to a hotspot that provides access to the local network where the server is running. In the experiment, an image is transmitted as a raw byte array over TCP, and the server performs inference using the received array to determine the predicted class, which is then returned as a numerical value. The duration from the start of the offloading process to the reception of the response is measured and averaged over $10,000$ images. Measurements for this process with both datasets are presented in Table V.
V-C Device Results
@@
The results for base models were acquired on 3 devices—Coral Micro, Raspberry Pi, and Jetson Orin-and are summarized in Table VI. TOn the Coral Micro, the results show that for CIFAR-10, AlexNet underperforms compared to ResNet models across all performance metrics. For ImageNet-1K, the trained AlexNet model could not be loaded onto the device due to its size, even after modifying the board’s memory regions via the linker script to allocate additional space. In contrast, all models were successfully executed on the Raspberry Pi. For Jetson Orin, an important observation is that operating at $15$W results in lower energy consumption compared to $7$W. This is because higher power enables lower latency, leading to greater overall efficiency. This phenomenon is similar to the trade-offs observed in BLE vs. Wi-Fi offloading times.
@@
For CIFAR-10, Jetson Orin is the most efficient device in terms of both energy consumption and latency compared to the other two devices. While the Raspberry Pi is less power-efficient than the Coral Micro for ResNet-8 ($\times$2.9) and ResNet-56 ($\times$5.6), an anomaly is observed with AlexNet, which is significantly more efficient on the Raspberry Pi ($\times$5.8) than on the Coral Micro.
@@
For ImageNet-1K, a similar pattern emerges. However, when considering latency, the Raspberry Pi outperforms the Coral Micro for ResNet-50 by a factor of 1.5. This suggests that running larger models on the Coral Micro incurs additional processing time, even with the assistance of its TPU accelerator.
VI On-Device vs. HI vs. Remote Inference
@@
In Fig. 4, we present the accuracy achieved by different strategies for CIFAR-10 and ImageNet-1K datasets, respectively. The corresponding latency and energy consumption on on Coral Micro, Raspberry Pi, and Jetson Orin in Figs. 5 and Fig. 6, respectively. ResNet-8 and ResNet-18, the state-of-the-art (SOTA) tinyML models for CIFAR-10 and ImageNet-1K, respectively, are used as the local DL models in the HI system. From Figs. 5 and 6, we observe similar trends across both datasets for Coral Micro and Raspberry Pi. In all cases, the smallest local models—ResNet-8 for CIFAR-10 and ResNet-18 for ImageNet-1K—consume the least energy and exhibit the lowest latency. However, these models also deliver the lowest accuracies (cf. Fig. 4) and fail to meet the accuracy QoS requirements.
(a) Accuracy on CIFAR-10
(b) Accuracy on ImageNet-1K
Figure 4: Accuracy of different strategies on CIFAR-10 and ImageNet-1K datasets
(a) Latency for CIFAR-10
(b) Energy consumption for CIFAR-10
@@
Figure 5: Latency and energy comparison of on-device and HI systems for CIFAR-10 on different devices.
(a) Latency for ImageNet-1K
(b) Energy consumption for ImageNet-1K
@@
Figure 6: Latency and energy comparison of on-device and HI systems for ImageNet-1K on different devices.
@@
The HI system achieves higher accuracy—up to $2.2$% higher for CIFAR-10 (cf. Fig. 4(a)) and $11.5$% for ImageNet-1K (cf. Fig. 4(b))—compared to both on-device models. This improvement comes with only a slight increase in latency and energy consumption relative to the smallest model (cf. Figs. 5 and 6). Remote ML inference yields the highest accuracy for CIFAR-10 but results in significantly higher latency and energy consumption on both Coral Micro and Raspberry Pi, as shown in Figs. 5(a) and 5(b). For ImageNet-1K, however, offloading data for remote ML inference leads to lower energy consumption and latency on non-GPU devices, due to the increased size of local ML models and the complexity of the dataset (cf. Figs. 6(a) and 6(b)). In contrast, Jetson Orin, with its GPU acceleration, is well-suited for on-device inference, as larger models like ResNet-56 for CIFAR-10 and ResNet-50 for ImageNet demonstrate significantly lower latency and energy consumption compared to remote ML inference (cf. Figs. 5, and 6).
@@
Next, we compare the performance of three systems—on-device inference, the HI system, and remote inference—based on the QoS requirements defined in Section III-C. For Arduino Nano and ESP32, in Fig. 1 shows that on-device inference using LR meets the accuracy QoS for the MNIST dataset, with significantly lower latencies compared to offloading. for CIFAR-10, the state-of-the-art tinyML model ResNet-8 fails to meet the $90$% QoS requirement, necessitating the use of a larger model. The high latencies observed for ResNet-8 on these MCUs indicate that larger models would result in even greater latency and energy consumption, making remote inference the preferred strategy for meeting QoS requirements.
@@
For Coral Micro and Raspberry Pi, we observe from Figs. 5(a) and 5(b) that for the CIFAR-10 dataset, both the on-device ResNet-56 model and the HI system with ResNet-8 can achieve the required accuracy QoS of $90$%. Although the HI system introduces some latency and energy overheads compared to the on-device ResNet-8 model, these are minimal when compared to the significantly higher latency and energy consumption required for on-device inference with ResNet-56. As a result, the HI system meets the accuracy QoS with up to $73$% lower latency and $77$% less energy than the on-device model. Given the latency and energy QoS requirements (cf. Table V), the HI system also achieves an absolute accuracy gain of over $2$%. For Jetson Orin, on-device inference with ResNet-56 is much faster and more energy-efficient than the HI system. Therefore, for the $90$% accuracy requirement, on-device inference is sufficient. However, since the HI system achieves higher accuracy ($1.64$% more) while still meeting the energy and time QoS requirements, it remains the best strategy in those scenarios.
TABLE VII: Preferred strategy for QoS requirements.
@@
Figure 7: EE-HI for a Convolutional Neural Network (CNN) with a single EE branch.
From Table VI and Fig. 4(b), we observe that for the ImageNet-1K dataset, the on-device DL models ResNet-18 and ResNet-50, with accuracies of $69.7$% and $76$% respectively, fall well short of the required accuracy QoS of $80$%. Furthermore, both models consume significantly more energy and exhibit higher latency on Coral Micro and Raspberry Pi compared to remote inference (cf. Figs. 6(a) and 6(b)). This makes using a larger local model on these two devices infeasible. In contrast, offloading data samples for remote inference achieves higher accuracy with significantly lower latency and energy consumption. Therefore, remote inference is the optimal strategy for meeting the $80$% accuracy QoS requirement. Additionally, it is the fastest and most energy-efficient solution for non-GPU devices.
On Jetson Orin, however, the on-device inference is faster and more energy-efficient. Therefore, we choose the RegNetY32GF model for on-device inference, as both ResNet-18 and ResNet-50 fail to meet the $80$% accuracy requirement. Using RegNetY32GF, we achieved $82.09$% accuracy with an inference time of $14.88$ms and energy consumption of $213.67$mJ, which is higher than that of remote inference. In contrast, the HI system with ResNet-18 meets the accuracy QoS with $50$% lower latency and $72$% lesser energy. Further, HI improves accuracy by $2.14$% over RegNetY32GF while satisfying both latency and energy QoS. The best strategies for each QoS requirement are summarized in Table VII.
@@
It is noteworthy that the decision module plays a crucial role in the success of HI systems, as demonstrated in (2). In our experiments, the binary LR performed well, achieving F1 scores of 0.86 on ResNet-8 for CIFAR-10 and 0.83 on ResNet-18 for the ImageNet-1K dataset. A key insight from this performance comparison is that:
@@
  • In general, on resource-constrained devices, relatively complex tasks—such as CIFAR-10 on the Arduino Nano and ESP32, or ImageNet-1K on the Coral Micro and Raspberry Pi—demand significantly higher energy and latency to meet QoS requirements with on-device inference. As a result, offloading these tasks for remote inference proves to be a more efficient choice.
@@
  • If on-device inference is feasible, it is often more advantageous to choose an HI system with a smaller model that does not meet the QoS requirement rather than a larger on-device model that does. HI systems typically satisfy QoS requirements while optimizing latency and energy consumption, balancing trade-offs between accuracy and resource efficiency.
VII Early Exit with HI
VII-A Algorithm Description
From (3), it can be inferred that the latency of an HI system depends on four factors: (i) on-device model latency, (ii) HI decision module latency (binay LR in this case), (iii) fraction of offloaded samples, and (iv) offloaded sample latency, which accounts for both the communication time and the computation time at the server. Thus, reducing any of these factors will reduce the overall latency per inference in the HI system. The same conclusion applies to the energy consumption of the HI system (cf. (4)). Recall that HI incurs latency and energy overhead due to the on-device model as every sample is first processed on it. To mitigate this overhead, we use the early exit technique [24], which has been proven to reduce on-device overhead while maintaining comparable accuracy. Specifically, we refer to the HI system using an on-device model (base model) with early exits as EE-HI system. An overview of an EE-HI system using CNNs is shown in Fig. 7.
We design and train the base models with early exit branches, referred to as EE models, following the pioneering BranchyNet approach [24]. Developing an EE model is a two-step process. First, an EE model with the desired architecture is created by adding branches with exit points between the base model’s layers. An EE model with a single exit branch at layer $k$ is shown in Fig. 7. Note that an EE model possesses one input and multiple outputs—one for each exit branch introduced during the design phase and the final layer. Then, the EE model is trained with a joint loss function, calculated as a weighted sum of the individual loss functions for each exit branch. Let $C$ denote the set of output labels, and $w_{n}$ denote the relative weight assigned to the exit branch indexed by $n$. We determine the weights $w_{n}$ following the guidelines provided in the BranchyNet approach [24]. Let $\textbf{y}=\{y_{c}:c\in C\}$ denote the ground truth vector, where $y_{c}\in\{0,1\}$ is $1$ for the ground truth label and 00, otherwise. The output of the EE model is denoted by $\hat{\textbf{y}}_{c}$, a concatenation of the vectors $\hat{\textbf{y}}_{\text{exit}_{n}}=\{\hat{y}_{nc}:c\in C\}$, where $\hat{y}_{nc}$ is the predicted (confidence) value for class $c$ at exit $n$. Each branch’s loss is defined as the Cross-Entropy loss between y and $\hat{\textbf{y}}_{\text{exit}_{n}}$. Thus, the joint loss function, denoted by $L_{\text{branchynet}}(\hat{\textbf{y}},\textbf{y})$, that we use to train the EE model, is given by
$L_{\text{branchynet}}(\hat{\textbf{y}},\textbf{y})=\sum_{n=1}^{N}w_{n}L(\hat{\textbf{y}}_{\text{exit}_{n}},\textbf{y}),$
where
$L(\hat{\textbf{y}}_{\text{exit}_{n}},\textbf{y})=-\frac{1}{|C|}\sum_{c\in C}y_{c}\log\hat{y}_{nc}.$
Several variables must be considered when designing an EE architecture, as they significantly impact model performance. The most critical factors are the number of branches ($N$), their placement, and their composition. The placement and composition of the EE branches are determined through trial-and-error, following the guidelines established in previous works [24, 46]. Additionally, for each branch, it is necessary to define the confidence threshold that determines whether to accept the inference at the given exit or continue processing in subsequent layers. We identify the optimal threshold using a brute-force approach. The optimal parameter values for each model are shown in Table VIII.
@@ VII-A Algorithm Description
For ResNet-8, we place one EE branch after the second residual block (or fifth convolutional layer), and for AlexNet, one EE branch is added after the second convolutional layer. For ResNet-56, two EE branches are used: the first is placed after layer $2$, and the second after layer $19$. Larger models can accommodate more branches without introducing excessive overhead relative to the base model’s latency. As noted earlier, the brute-force approach is used to compute the optimal threshold combination that satisfies the QoS requirements for each scenario.
TABLE VIII: Early Exit Parameters in EE-HI
Model EE Branch Placement Threshold ($\theta$)
ResNet-8 Layer 5 $\theta=0.69$
ResNet-56 1st: Layer 2, 2nd: Layer 19 $\theta_{1}=0.81,\theta_{2}=0.84$
AlexNet Layer 2 $\theta=0.75$
During the inference phase, once the model is successfully trained, a sample is first evaluated at the initial EE branch. If the confidence level (maximum softmax value) for classification exceeds the threshold, the system accepts the output and proceeds to the next sample. Otherwise, the sample continues to subsequent layers for further processing. This ensures that additional processing overhead is incurred only when necessary.
@@
TABLE IX: Results for the EE-HI system with optimal threshold subject to the QoS requirement.
@@
Figure 8: EE, HI, EE-HI comparison for different models for CIFAR-10 on Raspberry Pi.
VII-B Results
We evaluated the HI and EE-HI systems on Raspberry Pi and Jetson Orin using the CIFAR-10 dataset. The minimum accuracy QoS requirement for CIFAR-10 is 90%, and the results are summarized in Table IX. For each model and device, EE-HI consistently outperforms the base model with the HI strategy, achieving reductions of up to $60$% in both latency and energy consumption for ResNet-56 on Raspberry Pi.